40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6596     Entry time: Thu May 3 13:19:17 2012
Author: Jenne 
Type: Update 
Category: Locking 
Subject: More success last night 

I locked each arm, and the Michelson last night, no problems after I increased the Yarm gain from 0.1 to 0.2 .  I checked the green beam alignment just before going home, and both of the green beams are locking on ~03 or 04 modes, so aligning them is on the list for today.

ELOG V3.1.3-